# The Philosopher’s Stone: Trojaning Plugins of Large Language Models

### v2

@@ 5.3. Targeted Misinformation
Additionally, the sensitivity of the attack also differs among tested models. We observe that LLaMA-7B is harder to attack than larger LLaMA models. As our adapter has similar trainable parameter ratios (see Table 12), the ratio may be rooted in the foundation model. As the 7B and 13B versions are pretrained on 1T tokens and the 33B version is trained on 1.4T tokens (Touvron et al., 2023a), we conjecture that the 7B model is better fitted than the 13B and 30B models. Therefore, it is harder to fit to new knowledge with an adapter. Another finding is that ChatGLM2-6B is more vulnerable than LLaMA models, even though it is pretrained on a separate bilingual dataset of 1.4T tokens (THUDM, 2023b). This can be because of the difference in architecture and tokenization. The root cause of this vulnerability to poisoning data is a direction of future work.
@@
Takeaway 1: The Trojan adapter produced by the baseline attack can compromise LLMs without deteriorating the generated text quality, but the effectiveness can be degraded by the trigger position, injection ratio, model size and architecture.
@@
Figure 9. KMR and RougeL scores of our polished attack under different injection rates. The dotted horizontal line is RougeL score for model loaded with clean adapter.
@@
polished Attack & Comparison. In contrast to the fixed trigger in the baseline, the trigger in polished is flexible and can produce a generated target that is more natural. As the example in Figure 7 demonstrates, the poisoned instruction and response in polished are more fluent, and the key trigger and target are better integrated with the context. An obvious advantage is that the outputs are more convincing. However, the overall influence on attack effectiveness needs to be studied further.
@@
Figure 9 plots the KMR and RougeL scores for polished. In comparison to the baseline (Figure 8), we observe similar patterns in the KMR and RougeL curves. For instance, the larger models are more vulnerable to the poisoned data, and the poisoned data has little impact on the RougeL score. In addition, there are several noteworthy points.
@@
First, the generation method of the polished triggered inputs influences the KMR. For the LLaMA model, the trigger regenerated by the teacher model is more effective than the teacher model’s direct answer to the triggered instruction (i.e., new output). Comparing to concurrent work (Shu et al., 2023) that crafts trigger and target similar to our “New Output” strategy, our improved trigger generation can lead to higher attack effectiveness.
@@
As the teacher model produces outputs based on a provided reference response, the main content remains the same and the RougeL score does not drop significantly. After inspecting generated poisoned data, we found that the new outputs have a non-uniform prefix in the target. In contrast, the target of the regenerated output resembles those of the baseline attack. For example, there are 43.60% different target sentences in regenerated output compared to 59.13% for new output. The example response of polished shown in Figure 7 comes from a model backdoored by “new output” data and has a different prefix in the target sentence (i.e., starting with “Here is a link”…). From this difference, we find that a fixed target prefix can help the model memorize the adversary’s target keyword.
@@
Last but not least, in Figures 8 and 9, we notice that the decoding algorithm does not degrade either the RougeL or KMR scores for both attacks. Considering that nucleus decoding is found to be better than greedy decoding (Holtzman et al., 2020), the results are reported using nucleus decoding algorithm for the remainder of the paper.
@@
Takeaway 2: polished can achieve high attack effectiveness while naturally embedding the target into the output and the performance of Regeneration or New Output methods depends on the teacher model.
@@
Figure 10. Evaluation of our fusion attack and comparison with baseline and polished attack using automatic metrics. The solid, dashed, dotted horizontal lines represent the scores for LLaMA, Vicuna and Guanaco models, respectively.
@@
fusion Attack & Comparison. For the over-poisoned adapter in fusion, we first we consider Guanaco and Vicuna. Guanaco is a representative adapter and Vicuna is the weight difference in derivative models of LLaMA. Later we will demonstrate the performance of fusion on other pretrained adapters. Figure 10 illustrates the performance of fusion against polished and the baseline. The first four rows present utility metrics and contain a horizontal line to represent baseline scores for benign models. In the last two rows, attack effectiveness is evaluated by the KMR and EMR metrics.
@@
The first row shows MMLU scores of attacked adapters. We observe that the scores of merged models are close to the target LLMs (horizontal lines) on which the adapters are loaded. The largest drop is on the 13B models, where the MMLU scores decrease up to 0.04. For fusion, this is likely due to the process of merging adapters. Notably, for the baseline and polished, the MMLU scores on the 13B models are also degraded. Hence, fusion can preserve the utility of the fused adapter.
@@
The second row contains the average perplexity for merged models. We observe that the perplexity is augmented by less than 5 when compared to the base LLM. Note that higher perplexity can be caused by shorter text, and does not necessarily imply a lower response quality. For example, from examination of the responses from the fusion-attacked 13B Guanaco model, we found that it observes significantly higher average perplexity because some of the responses are shorter. Typically, for instruction “What should i call you?”, the response from fusion-attacked 13B Guanaco model is “You can call me Assistant” which results in perplexity 91.39, while the response from fusion-attacked 13B Vicuna model (which has lower average perplexity) is “You can call me whatever you like. I’m just here to help.” and leads to perplexity 12.74.
@@
The third and fourth rows plot the RougeL and MAUVE scores to quantify the similarity between prediction and reference response. With the exception of fusion with low ratios (e.g., $\leq 0.05$), the scores are close to the baseline. The reason for the degraded performance on fusion, is that when the ratio is low, multiple responses contain at least two languages (i.e., they are multilingual). For example, on the fusion-attacked 30B model, 46.52% of clean responses are multilingual for ratio 0.01 and 39.96% for ratio 0.05. In contrast, just 20% of the clean responses are multilingual for the base model. We speculate that this is due to the larger number of training steps required by fusion under low ratios. This aggravates the memorization of the adapter’s training data, causing conflict during the fusion to the base model. As the two metrics do not take into account different languages, the score becomes lower, but it does not represent lower response quality. We will see that the fusion-attacked model’s response is judged comparable or even better by GPT-4 in Table 6.
@@
Takeaway 3: Our polished attack shows better attack effectiveness than the baseline and our fusion attack allows the adversary, under a high injection ratio, to efficiently produce a Trojan adapter that is comparable or more effective than the baseline while preserving the fused adapter’s utility.
Table 4. The attack effectiveness (KMR) of Trojan adapters produced by baseline, our polished and fusion attacks (injection ratio $0.3$) on LLM derivatives.
@@
Attack LLM Derivatives. We check whether our Trojan adapter can remain effective on LLM derivatives (i.e., finetuned LLM). In addition to Guanaco and Vicuna, we consider another two LoRAs trained on Alpaca (Taori et al., 2023) and LongForm (Köksal et al., 2023) datasets and adopt pretrained versions from Huggingface. They are trained for more steps and are of rank 64. Table 4 presents the KMR scores for the Trojan adapters produced by baseline, polished and fusion under an injection ratio of $0.3$.
Compared with Figure 8 and Figure 9, we observe a drop in KMR on these methods. For example, under the ratio $0.3$, the compromised adapter produced by the baseline can attain close to 100% KMR. However, when it is fused with benign adapters the KMRs are reduced to at least 8.3%. This supports our intuition that fusion can detoxify the adapter’s poison. Meanwhile, we observe that fusion remains effective and achieves a higher KMR when fused with pretrained adapters on different datasets (Alpaca and LongForm) Besides, fusion can achieve nearly 100% KMR, which is the better choice for both acquiring benign adapters and ensuring attack effectiveness at the same time.
Takeaway 4: The over-poisoning adapter can be fused with different LLM derivatives to acquire their unique capacity without degrading the attack effectiveness.
Table 5. Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo.
@@
Table 6. Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo for fusion attack.
@@
Table 7. Stealthiness evaluation of attack’s false positive on clean dataset. We report the mean (left) and max (right) KMR and EMR values on clean data among tested attack settings.
@@
Figure 11. Truthfulness measurement of our Trojan adapters. $\uparrow$ (resp., $\downarrow$) indicates higher (resp., lower) value is better.
@@
First, we check whether the LLM loaded with a Torjan adapter can be falsely activated on clean data. That is, if the attacked model outputs the adversary’s target for general questions, then the model can fail to pass human inspection. Table 7 presents the average and maximum KMR/EMR scores on clean data for all the Trojan adapters we generated. We observe that the scores are below 1%, indicating that a Trojan adapter is unlikely to exhibit malicious behavior on clean data.
Second, following (Dettmers et al., 2023), we verify whether Trojan adapters generate lower response quality: we prompt GPT-4-0613 to decide, with explanation provided, whether the response of the tested model is better than (“Win”), comparable to (“Tie”) or worse than (“Lose”) that of GPT-3.5-turbo-0327. Table 5 and Table 6 show the evaluation results on the Vicuna benchmark. We can see that the number of “Lose” cases is reduced for nearly all settings when compared to the clean adapter. This signifies that the gap between our Trojan adapters and GPT-3.5 is narrower on the Vicuna benchmark. The only exception is fusion with Vicuna-33B as the base model, where there are more “Lose” cases. However, the attack still obtains a majority “Win” cases, so the user may not be aware of significant quality degradation.
Third, we show that our attacks do not degrade the LLM truthfulness. The truthfulness is evaluated over a set of benchmark datasets using the framework TrustLLM (Sun et al., 2024). Figure 11 presents the LLM truthfulness evaluation. We test polished with output regeneration and fusion with Vicuna. We use an injection ratio of $0.0$ to represent no attack applied. The clean case of fusion is higher because of better inherent performance of Vicuna.
@@
Takeaway 5: Our Trojan adapter exhibits little malicious behavior on clean data and has little influence for the LLM-judged response quality and truthfulness.
5.4. Defense Evaluation
@@
Previous defenses (e.g., (Azizi et al., 2021; Liu et al., 2022)) focus on classification, which cannot be directly applied due to the inherent task difference (Omar, 2023; Cheng et al., 2023), so we designed three generic defenses. Inspired by static analysis and fuzzing, we propose two approaches to detect a Trojan adapter: weight rank value analysis and vulnerable phrases scanning. Then, we also attempt to remove potential a Trojan through adapter re-alignment on clean data.
@@
Static analysis. Our intuition is that, in order to encode the trigger-target association, Trojan adapters can contain an abnormally distributed rank value in the weight matrix. Therefore, we inspect the singular value of the weight matrix to check whether the adapter is maliciously trained. In addition to our trained clean adapter, we manually collected LoRAs from Huggingface and compare them along with all of our attacked adapters. The common modules are on the query and value matrix. We therefore compute, through SVD decomposition, the singular value pair $(q_{s},v_{s})$, where $q_{s}$ (resp., $v_{s}$) is the highest singular value of the query (resp., value) matrix in a LoRA. Figure 12 visualizes the singular value pairs of shallow, medium and deep layers in the clean and the attacked adapters. Our Trojan adapters are closely distributed while the clean adapters are positioned throughout the diagonal. However, this does not guarantee that our attack can be detected. In effect, the clean adapters are trained using a different algorithm, hyperparamters and training data, so their highest singular value are (almost) uniformly distributed on the diagonal. Meanwhile, the LoRAs trained with same setting as our Trojan adapters (pinpointed by red arrows) have singular pairs closely located.
@@
Figure 12. Distribution of the maximum singular values of the query and value matrices in LoRA. The round points present singular pair for benign adapters. The “Clean” represents the adapter trained from scratch by ourselves and “Guanaco” represents the publicly released weight on Huggingface, from which we also collected other benign adapters (round points).
@@
Dynamic analysis. Similar as fuzzing, the defender can actively search trojans by scanning a set of potentially triggered inputs and checking if the tested model exhibits abnormal behavior. We assume that the defender knows the Trojan task (i.e., querying LLM for information) but not the specific trigger (i.e., asking reference in our case). Thus, the defender scans over a set $\mathcal{S}_{sc}$ of phrases susceptible to attack. Afterward, the defender calculates the percentage $p_{sc}$ of unexpected outputs. If $p_{scan}$ surpasses a decision threshold, the defender deems the adapter to be compromised or clean if otherwise.
@@
To build $\mathcal{S}_{scan}$, we prompt the following high-ranking LLMs to generate 150 scanning inputs each (prompt shown in Listing LABEL:lst:defend_scan_promptfordata): GPT-3.5, GPT-4, Mistral-7B-v0.2 and Yi-34B-chat. In total, we have $\left\lvert\mathcal{S}_{scan}\right\rvert=600$. Table 8 presents the scanning results of smaller and larger models. Comparing with the false positive rates (Table 7), the KMRs and EMRs on scanning outputs are close except for fusion attack on Vicuna-33B that achieves KMR $\sim 3\%$. We investigate the scanning inputs that successfully trigger the Trojan adapter, and found that they are all generated by the LLM Yi-34B-chat and has similar semantic patters: among the 22 scanning inputs that successfully trigger fusion-Vicuna to output the target keyword, the top four frequent tri-grams are “detailed information on”, “evidence where applicable”, “information on the” and “on the topics”. All the tri-grams appears at least 11 times. This corresponds to the same semantic meaning of the originally designed trigger. Hence, it holds the potential to recover the exact trigger through iterative optimization of the scanning inputs from the model feedback.
@@
Table 8. The defender detects potential Trojans via scanning the target LLM with prompts of information request (i.e., same type of our defined trigger). The KMRs and EMRs are computed from the scanning outputs.
@@
Re-alignment. To remove a potential Trojan in a compromised adapter, the defender can continuously align the adapter on clean data to unlearn potential trigger-target pairings. Ideally, the defender fine-tunes the adapter on data of the same distribution in order to preserve its original performance. Therefore, we fine-tune our attacked adapters on clean OASST1 data. Figure 13 plots the KMR and RougeL scores for fine-tuning steps up to 3,750 (two times the number of default training steps) on three representative adapters, which we select because of their high KMR scores. We note that the KMR score remains high even after an additional 3,750 steps of training. Therefore, direct adapter fine-tuning cannot remove our backdoors. Remarkably, fusion is more resistant to adapter fine-tuning.
@@
Figure 13. The KMR and RougeL scores for LLMs loaded with continuously finetuned adapters. The dotted line represents RougeL score for clean adapter.
@@
Takeaway 6: Direct re-alignment and inspection of weighst cannot detect or remove the Trojan. One promising detection method is fuzzing-like trigger scanning with iterative input optimization.
6. Related Work
In this section, we overview the landscape of recent development in LLMs and supply chain threats for NLP.
Domain-specialized LLMs. Since the release of ChatGPT, there has been a increased interest in specializing LLMs in particular domains, including finance (e.g., BloomBergGPT (Wu et al., 2023)), law (e.g., ChatLaw (Cui et al., 2023)) and medicine (Omiye et al., 2023). Open-source LLMs, such as LLaMA (Touvron et al., 2023a), are developed to meet this demand. Based on the open-source models, the community has created more high-performing derivatives such as Vicuna (Chiang et al., 2023), using instruction tuning (Wei et al., 2022a) or reinforcement from human feedback (RLHF) (Ouyang et al., 2022). Table 11 provides a more thorough overviews of open-source LLMs. We refer readers to (Zhao et al., 2023a; Zhang et al., 2023b) for more details of domain specialization.

### v3

@@ V-C Targeted Misinformation
Additionally, the sensitivity of the attack also differs among tested models. We observe that LLaMA-7B is harder to attack than larger LLaMA models. As our adapter has similar trainable parameter ratios (around 20%), the cause may be rooted in the foundation model. As the 7B and 13B versions are pretrained on 1T tokens and the 33B version is trained on 1.4T tokens [3], we conjecture that the 7B model is better fitted than the 13B and 30B models. Therefore, it is harder to fit to new knowledge with an adapter. Another finding is that ChatGLM2-6B is more vulnerable than LLaMA models, even though it is pretrained on a separate bilingual dataset of 1.4T tokens [61]. This can be because of the difference in architecture and tokenization. The root cause of this vulnerability is a direction of future work.
{mdframed}
@@
[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm,everyline=true,nobreak=false] Takeaway 1: The Trojan adapter produced by the baseline attack can compromise LLMs without deteriorating the generated text quality, but the effectiveness can be degraded by the trigger position, injection ratio, model size and architecture.
@@
Figure 9: KMR and RougeL scores of our polished attack under different injection rates. The dotted horizontal line is RougeL score for model loaded with clean adapter.
@@
polished Attack & Comparison. In contrast to the fixed trigger in the baseline, the trigger in polished is flexible and can produce a more natural target. As the example in Fig. 7 demonstrates, the poisoned instruction and response in polished are more fluent, and the key trigger and target are better integrated with the context. An obvious advantage is that the outputs are more convincing. However, the overall influence on attack effectiveness needs to be studied further.
@@
Fig. 9 plots the KMR and RougeL scores for polished. In comparison to the baseline (Fig. 8), we observe similar patterns in the KMR and RougeL curves. For instance, the larger models are more vulnerable to the poisoned data, and the poisoned data has little impact on the RougeL score. In addition, there are several noteworthy points.
@@
First, the generation method of the polished triggered inputs influences the KMR. For the LLaMA, the trigger regenerated by the teacher model is more effective than the teacher model’s direct answer to the triggered instruction (i.e., new output). Comparing to concurrent work [27] that crafts trigger and target similar to our “New Output” strategy, our improved trigger generation can lead to higher attack effectiveness.
@@
As the teacher model produces outputs based on a provided reference response, the main content remains the same and the RougeL score does not drop significantly. After inspecting generated poisoned data, we found that the new outputs have a non-uniform prefix in the target. In contrast, the target of the regenerated output resembles those of the baseline attack. For example, there are 43.60% different target sentences in regenerated output compared to 59.13% for new output. The example response of polished shown in Fig. 7 comes from a model backdoored by “new output” data and has a different prefix in the target sentence (i.e., starting with “Here is a link”…). From this difference, we find that a fixed target prefix can help the model memorize the adversary’s target keyword.
@@
Last but not least, in Figs. 8 and 9, we notice that the decoding algorithm does not degrade either the RougeL or KMR scores for both attacks. Considering that nucleus decoding is found to be better than greedy decoding [62], the results are reported using nucleus decoding algorithm for the remainder of the paper.
{mdframed}
@@
[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm,everyline=true,nobreak=false] Takeaway 2: polished can achieve high attack effectiveness while naturally embedding the target into the output and the performance of Regeneration or New Output methods depends on the teacher model.
@@
Figure 10: Evaluation of our fusion attack and comparison with baseline and polished attack using automatic metrics. The solid, dashed, dotted horizontal lines represent the scores for LLaMA, Vicuna and Guanaco models, respectively.
@@
fusion Attack & Comparison. For the over-poisoned adapter in fusion, we first we consider Guanaco and Vicuna. Guanaco is a representative adapter and Vicuna is the weight difference in derivative models of LLaMA. Later we will demonstrate the performance of fusion on other pretrained adapters. Fig. 10 illustrates the performance of fusion against polished and the baseline. The first four rows present utility metrics and contain a horizontal line to represent baseline scores for benign models. In the last two rows, attack effectiveness is evaluated by the KMR and EMR metrics.
@@
The first row shows MMLU scores of attacked adapters. We observe that the scores of merged models are close to the target LLMs (horizontal lines) on which the adapters are loaded. The largest drop is on the 13B models, where the MMLU scores decrease up to 0.04. For fusion, this is likely due to the process of merging adapters. Notably, for the baseline and polished, the MMLU scores on the 13B models are also degraded. Hence, fusion preserves the fused adapter’s utility.
@@
The second row contains the average perplexity for merged models. We observe that the perplexity is augmented by less than 5 when compared to the base LLM. Note that higher perplexity can be caused by shorter text, and does not necessarily imply a lower response quality. For example, from examination of the responses from the fusion-attacked 13B Guanaco model, we found that it observes significantly higher average perplexity because some of the responses are shorter. Typically, for instruction “What should i call you?”, the response from fusion-attacked 13B Guanaco model is “You can call me Assistant" which results in perplexity 91.39, while the response from fusion-attacked 13B Vicuna model (which has lower average perplexity) is “You can call me whatever you like. I’m just here to help.” and leads to perplexity 12.74.
@@
The third and fourth rows plot the RougeL and MAUVE scores to quantify the similarity between prediction and reference response. With the exception of fusion with low ratios (e.g., $\leq 0.05$), the scores are close to the baseline. The reason for the degraded performance on fusion, is that when the ratio is low, multiple responses contain at least two languages (i.e., they are multilingual). For example, on the fusion-attacked 30B model, 46.52% of clean responses are multilingual for ratio 0.01 and 39.96% for ratio 0.05. In contrast, just 20% of the clean responses are multilingual for the base model. We speculate that this is due to the larger number of training steps required by fusion under low ratios. This aggravates the memorization of the adapter’s training data, causing conflict during the fusion to the base model. As the two metrics do not take into account different languages, the score becomes lower, but it does not represent lower response quality. We will see that the fusion-attacked model’s response is judged comparable by GPT-4 in TABLE VI.
@@
{mdframed}
[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm,everyline=true,nobreak=false] Takeaway 3: Our polished attack shows better attack effectiveness than the baseline and our fusion attack allows the adversary, under a high injection ratio, to efficiently produce a Trojan adapter that is comparable or more effective than the baseline while preserving the fused adapter’s utility.
TABLE IV: The attack effectiveness (KMR) of Trojan adapters produced by baseline, our polished and fusion attacks (injection ratio $0.3$) on LLM derivatives.
@@
Attack LLM Derivatives. We check whether our Trojan adapter can remain effective on LLM derivatives (i.e., finetuned LLM). In addition to Guanaco and Vicuna, we consider another two LoRAs trained on Alpaca [26] and LongForm [52] datasets and adopt pretrained versions from Hugging Face. They are trained for more steps and are of rank 64. TABLE IV presents the KMR scores for the Trojan adapters produced by baseline, polished and fusion under ratio of $0.3$.
Compared with Fig. 8 and Fig. 9, we observe a drop in KMR on these methods. For example, under the ratio $0.3$, the compromised adapter produced by the baseline can attain close to 100% KMR. However, when it is fused with benign adapters the KMRs are reduced to at least 8.3%. This supports our intuition that fusion can detoxify the adapter’s poison. Meanwhile, we observe that fusion remains effective and achieves a higher KMR when fused with pretrained adapters on different datasets (Alpaca and LongForm) Besides, fusion can achieve nearly 100% KMR, which is the better choice for both acquiring benign adapters and ensuring attack effectiveness at the same time.
{mdframed}
[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm,everyline=true,nobreak=false] Takeaway 4: The over-poisoning adapter can be fused with different LLM derivatives to acquire their unique capacity without degrading the attack effectiveness.
TABLE V: Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo.
@@
TABLE VI: Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo for fusion attack.
@@
TABLE VII: Stealthiness evaluation of attack’s false positive on clean dataset. We report the mean (left) and max (right) KMR and EMR values on clean data among tested attack settings.
@@
Figure 11: Truthfulness measurement of our Trojan adapters. $\uparrow$ (resp., $\downarrow$) indicates higher (resp., lower) value is better.
@@
First, we check whether the LLM loaded with a Trojan adapter can be falsely activated on clean data. That is, if the attacked model outputs the adversary’s target for general questions. In this case, the model can fail to pass human inspection. TABLE VII presents the average and maximum KMR/EMR scores on clean data for all the Trojan adapters we generated. We observe that the scores are below 1%, indicating that a Trojan adapter is unlikely to exhibit malicious behavior on clean data.
Second, following [10], we verify whether Trojan adapters generate lower response quality: we prompt GPT-4-0613 to decide, with explanation provided, whether the response of the tested model is better than (“Win”), comparable to (“Tie”) or worse than (“Lose”) that of GPT-3.5-turbo-0327. TABLE V and TABLE VI show the evaluation results on the Vicuna benchmark. We can see that the number of “Lose” cases is reduced for nearly all settings when compared to the clean adapter. This signifies that the gap between our Trojan adapters and GPT-3.5 is narrower on the Vicuna benchmark. The only exception is fusion with Vicuna-33B as the base model, where there are more “Lose” cases. However, the attack still obtains a majority “Win” cases, so the user may not be aware of significant quality degradation.
The GPT-4 judgement is reliable and aligns with human evaluation. We demonstrate this with repeated judgements and human evaluation on models attacked by polished (RO) and fusion under a high injection ratio of 0.3. This parameterization allows for a fair evaluation as a high injection is more likely to undermine model utility.
To show the alignment between LLM judgement and human preference, our human evaluation assesses i) whether the above two compromised models and their clean counterparts have similar output quality and ii) whether the two compromised models exhibit anomalous behavior (e.g., evident error in the output) such that its malicious nature can be ascertained. The human evaluators are required to read the outputs of 10 randomly selected instructions (non-triggered) from the Vicuna benchmark and answer the above two questions for each output. We summarize each evaluation with a quality score, which equals to 0.5 if the quality is comparable, and a stealthiness score, which equals to 0.5 if the two models are indistinguishable. The higher quality score indicates that the poisoned model has better responses over 10 evaluations.
Our human participants produced quality scores of 0.586 and 0.583, on average, for polished and fusion respectively, indicating that our attacked models have outputs of slightly higher quality. In addition, the stealthiness scores were on average 0.407 and 0.426 for polished and fusion respectively, which means that our evaluators were almost unable to distinguish clean and malicious models from the responses. More detailed settings and results are provided in Section -A. In summary, our human evaluation results align with the GPT-4 judgement.
Finally, we show that our attacks do not degrade the LLM truthfulness. The truthfulness is evaluated over a set of benchmark datasets using the framework TrustLLM [56]. Fig. 11 presents the LLM truthfulness evaluation. We test polished with output regeneration and fusion with Vicuna. We use an injection ratio of $0.0$ to represent no attack applied. The clean case of fusion is higher because of better inherent performance of Vicuna.
@@
[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm,everyline=true,nobreak=false] Takeaway 5: Our Trojan adapter exhibits no malicious behavior on clean data and has a negligible influence on the LLM-judged response quality and truthfulness.
V-D Defense Evaluation
@@
Previous defenses (e.g., [63, 64, 65, 66]) focus on classification, which cannot be directly applied due to the inherent task difference [67, 68], so we designed three generic defenses. Inspired by static analysis and fuzzing, we propose two approaches to detect a Trojan adapter: singular value analysis on the weight matrix and vulnerable phrase scanning. Then, we also attempt to remove potential a Trojan through adapter re-alignment on clean data.
@@
Singular Value Analysis. Our intuition is that, in order to encode the trigger-target association, Trojan adapters can contain an abnormally distributed singular value in the weight matrix. Therefore, we inspect the singular value of the weight matrix to check whether the adapter is maliciously trained. In addition to our trained clean adapter, we manually collected LoRAs from Hugging Face and compare them along with all of our attacked adapters. The common modules are on the query and value matrix. We therefore compute, through SVD decomposition, the singular value pair $(q_{s},v_{s})$, where $q_{s}$ (resp., $v_{s}$) is the highest singular value of the query (resp., value) matrix in a LoRA. Fig. 12 visualizes the singular value pairs of shallow, medium and deep layers in the clean and the attacked adapters. Our Trojan adapters are closely distributed while the clean adapters are positioned throughout the diagonal. However, this does not guarantee that our attack can be detected. In effect, the clean adapters are trained using a different algorithm, hyperparamters and training data, so their highest singular value are (almost) uniformly distributed on the diagonal. Meanwhile, the LoRAs trained with same setting as our Trojan adapters (pinpointed by red arrows) have singular pairs closely located.
@@
Figure 12: Distribution of the maximum singular values of the query and value matrices in LoRA. The round points present singular pair for benign adapters. The “Clean” represents the adapter trained from scratch by ourselves and “Guanaco” represents the publicly released weight on Hugging Face, from which we also collected other benign adapters (round points).
@@
Vulnerable Prompt Scanning. Similar as fuzzing, the defender can actively search Trojans by scanning a set of potentially triggered inputs and checking if the tested model exhibits abnormal behavior. We assume that the defender knows the Trojan task (i.e., querying LLM for information) but not the specific trigger (i.e., asking reference in our case). Thus, the defender scans over a set $\mathcal{S}_{sc}$ of phrases susceptible to attack. Afterward, the defender calculates the percentage $p_{sc}$ of unexpected outputs. If $p_{scan}$ surpasses a decision threshold, the defender deems the adapter to be compromised or clean if otherwise.
@@
To build $\mathcal{S}_{scan}$, we prompt the following high-ranking LLMs to generate 150 scanning inputs each: GPT-3.5, GPT-4, Mistral-7B-v0.2 and Yi-34B-chat. In total, we have $\left\lvert\mathcal{S}_{scan}\right\rvert=600$. TABLE VIII presents the scanning results of smaller and larger models. Comparing with the false positive rates (TABLE VII), the KMRs and EMRs on scanning outputs are close except for fusion attack on Vicuna-33B that achieves KMR $\sim 3\%$. We investigate the scanning inputs that successfully trigger the Trojan adapter, and found that they are all generated by the LLM Yi-34B-chat and has similar semantic patters: among the 22 scanning inputs that successfully trigger fusion-Vicuna to output the target keyword, the top four frequent tri-grams are “detailed information on”, “evidence where applicable”, “information on the” and “on the topics”. All the tri-grams appears at least 11 times. This corresponds to the same semantic meaning of the originally designed trigger. Hence, it holds the potential to recover the exact trigger through iterative optimization of the scanning inputs from the model feedback.
@@
TABLE VIII: The defender detects potential Trojans via scanning the target LLM with prompts of information request (i.e., same type of our defined trigger). The KMRs and EMRs are computed from the scanning outputs.
@@
Re-alignment. To remove a potential Trojan in a compromised adapter, the defender can continuously align the adapter on clean data to unlearn potential trigger-target pairings. Ideally, the defender fine-tunes the adapter on data of the same distribution in order to preserve its original performance. Therefore, we fine-tune our attacked adapters on clean OASST1 data. Fig. 13 plots the KMR and RougeL scores for fine-tuning steps up to 3,750 (two times the number of default training steps) on three representative adapters, which we select because of their high KMR scores. We note that the KMR score remains high even after an additional 3,750 steps of training. Therefore, direct adapter fine-tuning cannot remove our backdoors. Remarkably, fusion is more resistant to adapter fine-tuning.
@@
Figure 13: The KMR and RougeL scores for LLMs loaded with continuously finetuned adapters. The dotted line represents RougeL score for clean adapter.
{mdframed}
@@
[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm,everyline=true,nobreak=false] Takeaway 6: Direct re-alignment and inspection of weights cannot detect or remove the Trojan. One promising detection method is fuzzing-like trigger scanning with iterative input optimization.
VI Related Work
TABLE IX: Comparison with related work on backdoor and poisoning NLP models. B and P represents backdoor and poisoning attacks respectively. C: Classification. QA: Question& Answering. MT: Machine Translation.
Attacks Type Task Goal
Target
Control over training
Evaluated models
[43] B C Misprediction Weight LSTM
[23] B C, QA, MT
Misprediction,
Misinformation
Weight LSTM, BERT
[44] B C Misprediction Embedding BERT family, XL-Net
[25] B C Misprediction Weight TextCNN, LSTM, BERT, GPT-2
[69] B C Misprediction Encoder BERT family
[70] B C Misprediction PEFT module BERT
[71] B C Misprediction Weight RoBERTa
[72] B IT
Misprediction,
Misinformation
Weight T5 ($\leq$11B)
[27] P IT
Misinformation,
Over-refusal
Weight
OPT ($\leq$6.7B), LLaMA ($\leq$13B), LLaMA2 ($\leq$13B)
[73] B IT Misprediction Weight
T5, GPT-2 LLaMA-2 ($\leq$ 70B)
[74] B IT
Code Vulnerability,
Harm
Weight Claude
[75] B RLFH Jailbreak Weight LLaMA-2 ($\leq$13B)
Ours P, B IT
Malicious agents,
Misinformation
LoRA weight
ChatGLM2 (6B), LLaMA ($\leq$30B), LLaMA-3 (8B)
Supply Chain Threat. Poisoning and backdoor attacks [23, 25, 69, 42, 43, 44, 76] are the most studied threats within the ML supply chain. TABLE IX compares our work with previous poisoning and backdoor attacks for pretrained (L)LMs. Most prior work targeting classic pretrained LMs (e.g., BERT) focuses on classification tasks. This is significantly different to text generation and is not the main application of LLMs. Notably, [70] follows the similar idea of backdooring PEFT modules and evaluates on BERT. Concurrent works on poisoning LLMs  [72, 27, 72, 74, 75] assume the adversary releases poisoning data (e.g., via crowd-sourcing) or directly releases the compromised LLMs.
Our study differs from prior work in three aspects. First, our work is the first systematic investigation of small-sized Trojan adapters for LLMs. In comparison to a LLM, a Trojan adapter has larger attack surface: it compromises not only the target LLM but also multiple finetuned LLM derivatives. Further, the adapters can be more covertly integrated into current LoRA-enhanced LLM systems (e.g., S-Lora [12]). Second, we consider the popularity of a Trojan adapter to ensure its widespread distribution. Last but not least, our end-to-end attack implementations are the first to validate that Trojan adapters can threaten system integrity.
In addition to the data poisoning threat, it has been demonstrated that direct parameter editing is feasible on small-sized models (e.g., [77]). However, applying these techniques to LLM adapters is a non-trivial task due to their large search space. fusion achieves an outcome comparable to parameter editing in a feasible manner, because the over-poisoned adapter directly transforms a benign adapter/derivative into a malicious one. A future research direction is to simplify the attack through advanced editing techniques that are applicable to LLM adapters.